find the equation of the inverse of the following function
f(x)=6/x+7
y=6/x+7
y=7x+6/x
xy-7x=6
x(y-7)=6
x=6/y-7
f(y)=6/y-7
f(x)=6/x-7

Answer by solver91311(24713)


Your work:

f(x)=6/x+7

y=6/x+7

y=7x+6/x

xy-7x=6

x(y-7)=6







Yes, that is correct if the actual starting function was . However, if the original function was really , that is a horse of a completly different color.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism